OpenStudy (anonymous):

8e+3(5-e)=10 help, i dont get how to solve these

OpenStudy (anonymous):

are u trying to solve for e?

OpenStudy (anonymous):

8e+15-3e=10 8e-3e+15=10 5e+15-15=10-15 5e=-5 e=-1

OpenStudy (anonymous):

yes, solve for e